OpenStudy (anonymous):

In 1995 the life expectancy in a certain country was 71.4 years.In 2000,it was 74.8 years.Let E represent the life expectancy in the year t and let t represent the number of yers since 1995.The linear function E(t) that fits the data is E(t)=?t+?. Use the function to predict the life expectancy of males in 2004 E(9)=?

OpenStudy (saifoo.khan):

E(t)=0.68t+71.4

OpenStudy (saifoo.khan):

E(t)=0.68t+71.4 E(9)= 77.52

OpenStudy (anonymous):

thank you.

OpenStudy (saifoo.khan):

No problem! ^_^

OpenStudy (saifoo.khan):

Got more?

OpenStudy (anonymous):

That's all I've got right now. Thanks.

OpenStudy (saifoo.khan):

you welcome! :)

OpenStudy (anonymous):

Oh Sally and Bills ages total 28.Bill is 20 years older than Sally, how old is Sally?

OpenStudy (saifoo.khan):

x = bills y = sally x+y = 28 y +20 = x

OpenStudy (saifoo.khan):

Solving them, u will get bill = 24 sally = 4

OpenStudy (anonymous):

great answer and your an atheist.

OpenStudy (saifoo.khan):

Thanks. Thanks Alot. :)

OpenStudy (anonymous):

I'm right though aren't I.

OpenStudy (saifoo.khan):

are u talking about being an atheist?

OpenStudy (anonymous):

At least the Harvard research study concluded that, the people that figured out the Bill and Sally question usually were atheists.

OpenStudy (saifoo.khan):

lol, cool. never heard of that.. but im not an Atheist. hehe.

OpenStudy (saifoo.khan):

are u?